The probability of success on any trail of a binomial experiment is 25%. Find the probability that the proportion of successes in a sample of 1000 is between 22% and 28%.


1
Expert's answer
2022-05-24T13:18:38-0400

The probability of success for any trail of a binomial experiment is, p=0.25.

The sample size, n=1000

The probability of success that the proportion of success in a sample of 1000 is between 22% and 28% is,

"P(0.22<\bar{P}<0.28)=P( \frac{0.22-0.25}{\sqrt{\frac{0.25(1-0.25)}{1000}}} <\frac{\bar{P}-p}{\sqrt{\frac{p(1-p)}{n}}}<\frac{0.28-0.25}{\sqrt{\frac{0.25(1-0.25)}{1000}}} ) \\

=P(\frac{0.22-0.25}{0.0137}<Z<\frac{0.28-0.25}{0.0137}) \\

=P(-2.19<Z<2.19) \\
=\Phi(2.19)-\Phi(-2.19)\\
= 0.9857-0.0143=0.9714"

We found P using z-score table.